OpenStudy (anonymous):

how to find the area between two curves y=e^x, y=x-1, -2<=x<=0

OpenStudy (phi):

do you know calculus? the "height" of the region is the top curve minus the lower curve the width of a small rectangle is dx integrate along x from -2 to 0: height * dx see graph.

OpenStudy (anonymous):

i found it on the calculator but want to know it alegbraic

OpenStudy (phi):

what did you get for the area ? it's not algebra, it is calculus that you use.

OpenStudy (anonymous):

\[\int\limits_{-2}^{0}e ^{x} and \int\limits_{-2}^{0}x-1 \] i got 4.8646 on the calculator but i want to know how to get the answer w/o one

OpenStudy (phi):

the area is height * dx where height is the top curve e^x minus the lower curve x-1 i.e. height = e^x - (x-1) or height= e^x -x +1 and and the area is \[ \int_{-2}^0 e^x -x +1 \ dx \]

OpenStudy (phi):

You can "cheat" by noticing the area below the x-axis is a trapezoid (or a triangle on top of a rectangle), and has area 4 so we could say the total area is the area beneath the e^x curve plus the 4 \[ area = 4 + \int_{-2}^0 e^x \ dx \] this should also give the same answer.

OpenStudy (phi):

\[ \int_{-2}^0 e^x -x +1 \ dx \\ = e^x - \frac{x^2}{2} +x \ \bigg|_{-2}^0 \\ = (1 - 0 +0) - (e^{-2} -\frac{4}{2} -2) \\ = 5 - e^{-2} \] at this point, you need a calculator to evaluate e^-2 but if you don't have one, you can roughly estimate the expression: \( e \approx 2.7\). If we call it 3, then 3^2 = 9, and 3^-2 is 1/9 or about .11 so (very roughly) the answer is 5-0.11 = 4.89